OpenStudy (anonymous):

I'm going to cry help :c Solve each of the quadratic equations below and describe what the solution(s) represent to the graph of each. Show your work to receive full credit. y = x2 + 3x + 2 y = x2 + 2x + 1

OpenStudy (anonymous):

no crying over math

OpenStudy (anonymous):

Math is hard.

OpenStudy (anonymous):

yes that is true

OpenStudy (anonymous):

here is a graph of the first one http://www.wolframalpha.com/input/?i=y+%3D+x^2+%2B+3x+%2B+2+

OpenStudy (anonymous):

you can see that \[y=x^2 + 3x + 2 \] factors as \[y=(x+2)(x+1)\]

OpenStudy (anonymous):

this means if you want to find out where the graph crosses the \(x\) axis it is relatively easy. it crosses the \(x\) axis where \(y=0\) so set \[(x+2)(x+1)=0\] and solve for \(x\)

OpenStudy (anonymous):

you get \(x+2=0\implies x=-2\) or \(x+1=0\implies x=-1\) so you know where it crosses the \(x\) axis at the points \((-2,0)\) and \((-1,0)\) you can see it from the picture i sent

OpenStudy (anonymous):

does this make any sense to you?

OpenStudy (anonymous):

No LOOOL :c

OpenStudy (anonymous):

any question i can answer? that was my best explanation also a nice picture

OpenStudy (anonymous):

Can you show me how to like factor it out? :c Sorry I didn't take algebra last year so that's why I'm all confused on it.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

we start with \[y=x^2+3x+2\] and we want to make it look like \[y=(x+a)(x+b)\]

OpenStudy (anonymous):

so we need two numbers, \(a\) and \(b\) with \(a\times b=2\) and also \(a+b=3\)

OpenStudy (anonymous):

it is pretty clear that the two numbers that work are 1 and 2 because \(1\times 2=2\) and \(1+2=3\)

OpenStudy (anonymous):

so we can see that \[x^2+3x+2=(x+1)(x+2)\] and we can check that it is correct by multiplying out on the right and see that we get what we want

OpenStudy (anonymous):

that is, if you multiply \[(x+1)(x+2)\] you get \[x^2+x+2x+2=x^2+3x+2\] so we know it is right

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

OH!! Thanks so much ^__^

OpenStudy (anonymous):

Yeah I get it c:

OpenStudy (anonymous):

good can you do the next one?